The solubility of an unknown compound, c, in ethanol is 24mg per 100ml at 20 celsius and 0.78g pee 50ml at 90 Celsius

A) calculate the minimum amount of ethanol needed to recrustallize 2.4g of compound c.

B) calculate the amount of compound c recovered by filtration at 20 Celsius from this minimum amount of ethanol.
Answer #1

A) The recrystallization will be done at boiling point of ethanol. So, we will consider solubility at 90 deg C. At this temperature, 0.78 g of compound C dissolves in 50 ml ethanol. So, volume of ethanol needed to dissolve 2.4 g compound C is:

(50 ml/0.78 g)*2.4 g = 153.8 ml

Volume of ethanol required = 153.8 ml

B) Amount of compound C left by filtration is:

(24 mg/100 ml)*153.8 ml = 36.9 mg = 0.037 g

Amount of compound C recovered after filtration = 0.780 - 0.037 g = 0.743 g